There are 140 students in 9th grade and 140 students in 10th grade at Kennedy High School. We know that 3/5 of the 9th graders and 3/4 of the 10th graders went to the basketball game. How many more 10th graders than 9th graders went to the game?
Answer by Maths68(1474)

9th grade students = 140
9th graders went to basketball game = 140*(3/5)
9th graders went to basketball game = (140*3)/5
9th graders went to basketball game = 420/5
9th graders went to basketball game = 84
10th grade students = 140
10th graders went to basketball game = 140*(3/4)
10th graders went to basketball game = (140*3)/4)
10th graders went to basketball game = 420/4
10th graders went to basketball game = 105
Total number of 10th graders than 9th graders went to the game = (10th graders went to basketball game) – (9th graders went to basketball game)
Total number of 10th graders than 9th graders went to the game = 105-84
Total number of 10th graders than 9th graders went to the game = 21
21 more 10th graders than 9th graders went to the game.